OpenStudy (anonymous):

PLEASE HELP!! What value of m will make this expression equal to 6.35? When m = -12.7, the expression will equal 6.35. Any value of m, except when m = -12.7, will make the expression equal to 6.35. m = 0 is the only value that will make the expression equal to 6.35. There is no value that will make the expression equal to 6.35.

OpenStudy (danjs):

im guessing it is a line y = m*x + b 6.35 = m*x + b need more info, like a point on the line, or the expression from the previous part of th eprob

OpenStudy (danjs):

or actually 3 will work, with no info,

OpenStudy (danjs):

if b=6.35

OpenStudy (anonymous):

thanks!

OpenStudy (danjs):

lol welcome, not sure if i got it, was guessing

OpenStudy (danjs):

assuming it is a line